A car negotiate a bend of radius 20m with an acceleration of 12m/s². What is the maximum speed of the car attained without skipping?
1
Expert's answer
2020-04-01T09:59:32-0400

Solution:


"a= \frac {v^2}{R}"

"v^2=aR"


"v= \sqrt {aR}"

"v= \sqrt {12\times 20}\approx19 \frac {m}{s}"

Answer: "\approx 19 \frac {m}{s}"